OpenStudy (anonymous):

A word problem: A jazz concert brought in $207,000 on the sale on 8,000 tickets. If the tickets sold for $20 and $30 each, how many of each type of ticket were sold?

OpenStudy (lilg132):

x + y = 8000 y = 8000 -x 20x + 30(8000 - x) = 207,000 20x + 240,000 - 30x = 207,000 20x - 30x = 207,000 - 240,000 -10x = -33000 10x = 33,000 10x / 10 = 33,000 / 10 x = 3300 8000 - 3300 = y y = 4700 dont forget to click good answer :)

hero (hero):

I did that just to prove to you that I'm not a hater

OpenStudy (lilg132):

i didnt say your a hater look i cant be bothered just forget what happened and leave it as that im off anyways bye

hero (hero):

You seem to have anger issues or something. What is wrong with you?

OpenStudy (lilg132):

have you started again?

OpenStudy (lilg132):

hero are you mentally retarded or something when i said forget about what i said its my of saying forgive and forget but you just seem to WANT to drag this on anyways if thats what you want i dont have time for that so i will just ignore you from now on